OpenStudy (anonymous):

A triangle has 3 sides that measure 5 1/4ft, 2 5/6ft, and 8 7/12ft. What is the preimeter of the triangle.

OpenStudy (anonymous):

1. 2/18 + 16/17 = (34 + 288)/306 = 322/306 = 1 and 8/153 2. 11/4 + 11/3 = 77/12 = 6 5/12 3. 11/3 * 33/4 = 363/12 = 30 1/4 4. 1/2-1/3 = 1/6 5. 2/3+3/2 = 13/6 = 2 1/6 6. 21/4+17/6+103/12 = (63+34+103)/12 = 200/12 = 16 2/3 7. 19/3-7/3 = 12/3 = 4 8. 2/3 * 12 = 24/3 = /8 9. 1/2 * 1/8 = 1/16 10. 21/4 * 3/2 = 63/8 = 7 7/8 11. (5/3)/(15/7) = (5/3) * (7/15) = 7/9 12. (3 5/6)/(1/3) = 23/6 * 3/1 = 23/2 = 11 1/2 pieces of 1/3-yd can be cut

OpenStudy (anonymous):

its 16 1/12ft or 15 2/3ft or 15 13/22ft or 16 2/3ft

OpenStudy (anonymous):

?